OpenStudy (anonymous):

Write the slope-intercept form of the equation of (-5,7) and (4,-2)

OpenStudy (anonymous):

Find slope first. Use the equation \[(y-y _{0})/ (x-x _{0})\] Then, take either point and plug it into y=mx + b, where m is the slope you just found. Find out what b is. Then re-write y=mx+b ONLY USING m and b. Leave x and y as x and y.

OpenStudy (anonymous):

y=11/13x+b.

OpenStudy (anonymous):

It is actually y=-x + 2, you can check this by substituting the values in the given....